# Harmonics of Learning: Universal Fourier Features Emerge in Invariant Networks

### v2

@@ Abstract
Figure 1: Emergent Circular Harmonics. Weights learned by a neural network trained for invariance to planar rotations resemble circular harmonics. Data from [44].
@@ 1 Introduction and Related Work
Artificial neural networks trained on natural data exhibit a striking phenomenon: regardless of exact initialization, dataset, or training objective, models trained on the same data domain frequently converge to similar learned representations [29]. For example, the early layer weights of diverse image models tend to converge to Gabor filters and color-contrast detectors [35]. Remarkably, many of these same features are observed in the visual cortex [20, 16, 50], suggesting a form of representational universality that transcends biological and artificial substrates. While such findings are empirically well-established in the mechanistic interpretability literature [41], the field lacks theoretical explanations.
@@
Spatially localized versions of canonical 2D Fourier basis functions, such as Gabor filters or wavelets, are perhaps the most frequently observed universal features in image models. They commonly arise in the early layers of vision models – trained with efficient coding [36, 4], classification [35], temporal coherence [23], and next-step prediction [11] objectives – as well as in the primary visual cortices of diverse mammals – including cats [21], monkeys [22], and mice [10]. Non-localized Fourier features have been observed in networks trained to solve tasks that permit cyclic wraparound – for example, modular arithmetic [34], more general group compositions [7], or invariance to the group of cyclic translations [44]. In the domain of spatial navigation, the so-called grid cells of the entorhinal cortex [33] display periodic firing patterns at different spatial frequencies as they build a map of space. Their response properties are naturally modeled with the harmonics of the twisted torus [15, 37, 13]. Similar features also emerge in artificial neural networks trained to solve spatial navigation tasks [2, 8, 47]. The ubiquity of these features across diverse learning systems is both striking and unexplained.
@@
In this work, we provide a mathematical explanation for the emergence of Fourier features in learning systems such as neural networks. We argue that the mechanism responsible for this emergence is the downstream invariance of the learner to the action of a group of symmetries (e.g. planar translation or rotation). Since natural data typically possess symmetries, invariance is a fundamental bias that is injected both implicitly and sometimes explicitly into learning systems [6]. Motivated by this, we derive theoretical guarantees for the presence of Fourier features in invariant learners that apply to a broad class of machine learning models.
@@
Our results rely on the inextricable link between harmonic analysis and group theory [12]. The standard discrete Fourier transform is a special case of more general Fourier transforms on groups, which can be defined by replacing the standard basis of harmonics by irreducible unitary group representations. The latter are equivalent to the familiar definition for cyclic or, more generally, commutative groups, but are more involved for non-commutative ones. In order to accommodate both the scenarios, we develop a general theory that applies, in principle, to arbitrary finite groups.
@@
This work represents an attempt to provide mathematical grounding for a general algebraic theory of representation learning, while addressing the universality hypothesis for neural networks [35, 32]. A suite of earlier theoretical works [24, 17, 14] established such universality for sparse coding models [36], deriving the conditions under which a network will recover the original bases that generate data through sparse linear combinations. In this case, the statistics of the data determine the uniqueness of the representation. Our findings, on the other hand, are purely algebraic, since they rely exclusively on the invariance properties of the learner. Given the centrality of invariance to many machine learning tasks, our theory encompasses a broad class of scenarios and neural network architectures, while providing a new perspective on classical neuroscience [39, 21]. As such, it sets a foundation for a learning theory of representations in artificial and biological neural systems, grounded in the mathematics of symmetry.
@@ 1.1 Overview of Results
In the above, the term ‘harmonic’ refers to an irreducible unitary representation of $G$. Indeed, one-dimensional unitary representations correspond to homomorphisms with the unit circle $\textnormal{U}(1)\subseteq\mathbb{C}$, which is reminiscent of the classical definition via the imaginary exponential. However, non-commutative groups can have higher-dimensional irreducible representations, intuitively meaning that harmonics are valued in unitary matrices. In this case, the components of $W$ can be interpreted as capsules in the sense of [43] i.e., neural units processing matrix-valued signals.
We show that the hypothesis on $\varphi$ in Theorem 1.1 is satisfied by several machine learning models from the literature. In particular, the theorem applies to the recently-introduced (Bi)Spectral Networks [44], to single fully-connected layers of McCulloch-Pitts neurons, and, to an extent, to traditional deep networks. Since harmonic analysis is naturally formalized over the complex numbers, we consider models with complex weights $W$, which fits into the larger program of complex-valued machine learning [3, 49, 30].
The group-theoretical Fourier transform encodes the entire group structure of $G$. Therefore, as a consequence of Theorem 1.1 the multiplication table of $G$ can be recovered from the weights $W$ of an invariant parametric function $\varphi$ – a fact empirically demonstrated in [44]. This addresses the question of symmetry discovery – an established machine learning problem aiming to recover the unknown group of symmetries of data with minimal supervision and prior knowledge [40, 46, 9].
Since the multiplication table is a discrete object, it is expected that the invariance constraint on $\varphi$ can be loosened while still recovering the group correctly. To this end, we prove the following.
Informal Theorem 1.2 (Theorem 3.6).
If $\varphi(W,x)$ is ‘almost invariant’ to $G$ according to certain functional bounds and the weights are ‘almost orthonormal’, then the multiplicative table of $G$ can be recovered from $W$.
Lastly, we implement a model satisfying the requirements of our theory and demonstrate its symmetry discovery capabilities. To this end, we train it via contrastive learning on an objective encouraging invariance and extract the multiplicative table of $G$ from its weights. Our Python implementation is available at a public repository  [^1] https://github.com/sophiaas/spectral-universality.
@@ 2 Mathematical Background
We begin by introducing the fundamental concepts from harmonic analysis and group theory used in this paper. For a complete treatment, we refer the reader to [12].
@@ 2.1 Groups and Actions
Definition 1.
A group is a set $G$ equipped with a composition map $G\times G\rightarrow G$ denoted by $(g,h)\mapsto gh$, an inversion map $G\rightarrow G$ denoted by $g\mapsto g^{-1}$, and a distinguished identity element $1\in G$ such that for all $g,h,k\in G$:
@@
Examples of groups include the permutations of a set and the general linear group $\textnormal{GL}(V)$ of invertible operators over a vector space $V$, both equipped with the usual composition and inversion of functions. A further example that will be relevant in this work is the unitary group $\textnormal{U}(V)\subseteq\textnormal{GL}(V)$ associated to a Hilbert space $V$, consisting of operators $U$ satisfying $UU^{\dagger}=I$, where $\dagger$ denotes the conjugate transpose and $I$ is the identity matrix. Groups satisfying $gh=hg$ for all $g,h\in G$ are deemed commutative.
The idea of a space $\mathcal{X}$ having $G$ as a group of symmetries is abstracted by the notion of group action.
Definition 2.
@@ 2.2 Harmonic Analysis on Groups
Harmonic analysis on groups [42] generalizes standard harmonic analysis. We focus here on finite groups for simplicity, which are sufficient for practical applications. This avoids technicalities such as integrability conditions and continuity issues arising for infinite groups. We start by considering commutative groups and cover non-commutative ones in Section 2.3.
Let $G$ be a finite commutative group of order $|G|$. Denote by $\langle G\rangle=\mathbb{C}^{G}$ the free complex vector space generated by $G$. Intuitively, an element $x=(x_{g})_{g\in G}\in\langle G\rangle$ represents a complex-valued signal over $G$. The space $\langle G\rangle$ is endowed with the convolution product,
@@
Definition 3.
The dual $G^{\vee}$ of $G$ is the set of homomorphisms $\rho:\ G\rightarrow\textnormal{U}(1)$, where $\textnormal{U}(1)\subseteq\mathbb{C}$ is the group of unitary complex numbers equipped with multiplication. It is itself a group when equipped with pointwise composition $(\rho\mu)(g)=\rho(g)\mu(g)$.
A homomorphism $\rho\in G^{\vee}$ intuitively represents a harmonic over $G$, generalizing the familiar notion from signal processing. If we endow $\langle G\rangle$ with the canonical scalar product $\langle x,y\rangle=\sum_{g\in G}\overline{x_{g}}y_{g}$, then $G^{\vee}\subseteq\langle G\rangle$ forms an orthogonal basis with all the norms equal to $|G|$. The linear base-change is, by definition, the Fourier transform over $\langle G\rangle$:
Definition 4.
The Fourier transform is the map $\langle G\rangle\rightarrow\langle G^{\vee}\rangle$, $x\mapsto\hat{x}$, defined for $\rho\in G^{\vee}$ as:
$\hat{x}_{\rho}=\langle\rho,x\rangle.$ (2)
The Fourier transform is a linear isometry or, equivalently, a unitary operator, up to a multiplicative constant of $|G|$. Moreover, it exchanges the convolution product $\star$ over $\langle G\rangle$ with the Hadamard product $\odot$ over $\langle G^{\vee}\rangle$. Definition 4 generalizes the usual discrete Fourier transform in the following sense. For an integer $d>0$, consider the cyclc group $C_{d}$ with $d$ elements. Concretely $G=\mathbb{Z}/d\mathbb{Z}\simeq C_{d}$ is the group of integers modulo $d$ equipped with addition as composition. The dual $G^{\vee}$ consists of homomorphisms of the form:
$\mathbb{Z}/d\mathbb{Z}\ni g\mapsto e^{2\pi\sqrt{-1}gk/d},$ (3)
for $k\in\{0,\cdots,d-1\}$. Equation 2 specializes then to the familiar definition of the Fourier transform.
@@ 2.3 Non-Commutative Harmonic Analysis
So far, we have assumed that $G$ is commutative. In this section we briefly discuss the extension of Fourier theory to non-commutative groups. This however requires more elaborate theoretical tools, which we now introduce. To begin with, in order to perform harmonic analysis on general groups it is necessary to discuss unitary representations. The latter will play the role of matrix-valued harmonics.
Definition 5.
@@
This generalizes Definition 4 since for a commutative group, $\rho_{V}$ is irreducible if, and only if, $\textnormal{dim}(V)=1$. Analogously to the commutative setting, the Fourier transform exchanges the convolution product $\star$ with the point-wise operator composition, which we still denote by $\odot$. Moreover, the Fourier transform is a unitary operator up to a multiplicative constant of $|G|$ with respect to the normalized Hilbert-Schmidt scalar product on $\textnormal{End}(V)$, given by:
$\langle A,B\rangle=\textnormal{dim}(V)\ \textnormal{tr}(A^{\dagger}B).$ (5)
The norm associated to the Hilbert-Schmidt scalar product is the Frobenius norm. The relations between irreducible unitary representations coming from the unitarity of the Fourier transform are known as Schur orthogonality relations.
@@ 3 Theoretical Results
Let $G$ be a finite group, $\mathcal{H}$ be a set, and $V_{1},\cdots,V_{k}$ be complex finite-dimensional Hilbert spaces. In what follows, we will consider the space
@@
$\mathcal{W}$ is a Hilbert space when endowed with the scalar product given by the product of the canonical scalar product over $\langle G\rangle$ and the normalized Hilbert-Schmidt scalar products over $\textnormal{End}(V_{i})$ (see Equation 5). For $W\in\mathcal{W}$, we will denote each of its components as $W_{i}=(W_{i}(g))_{g}\in\textnormal{End}(V_{i})^{\oplus G}$. Moreover, we will often interpret elements $W\in\mathcal{W}$ as linear maps $\langle G\rangle\rightarrow\bigoplus_{i}\textnormal{End}(V_{i})$ via $W(x)=\sum_{g\in G}W(g)x_{g}$ for $x\in\langle G\rangle$, where $W(g)=(W_{i}(g))_{i}$. Note that $G$ acts on the left tensor factor of $\mathcal{W}$ while for each $i$, $\textnormal{U}(V_{i})$ acts on the right tensor factor of $\langle G\rangle\otimes\textnormal{End}(V_{i})$ by composition of operators.
Definition 7.
We say that a map $\varphi\colon\mathcal{W}\rightarrow\mathcal{H}$ has unitary symmetries if for $W,W^{\prime}\in\mathcal{W}$ of the same norm, $\varphi(W)=\varphi(W^{\prime})$ implies that for each $i$ there exists a unitary operator $U_{i}\in\textnormal{U}(V_{i})$ such that $W_{i}=U_{i}\cdot W_{i}^{\prime}$.
In the context of machine learning, $\mathcal{H}$ will represent the hypothesis space, consisting of functions the model can learn. On the other hand, $\varphi$ will represent the parametrization of such hypotheses, with its domain $\mathcal{W}$ being the space of weights. The component $\langle G\rangle$ of $\mathcal{W}$ will be responsible for parametrizing the input space, while each component $\textnormal{End}(V_{i})$ will represent a computational unit, i.e. a complex-valued neuron in the language of neural networks. For commutative groups, we simply have $V_{i}=\mathbb{C}\simeq\textnormal{End}(V_{i})$. In general, $\textnormal{End}(V_{i})$ can be thought of as parametrizing matrix-valued signals, which, as mentioned in Section 1.1, are computed by neural units sometimes referred to as capsules [43].
The following is an abstract algebraic principle at the core of this work.
@@
  • $W$, seen as a linear map $\langle G\rangle\rightarrow\bigoplus_{i}\textnormal{End}(V_{i})$, is surjective.
@@
Proof.
Since $\varphi$ has unitary symmetries and $\|g\cdot W\|=\|W\|$, it follows that for every $g\in G$ and every $i$ there exists $\rho_{i}(g)\in\textnormal{U}(V_{i})$ such that
$g\cdot W_{i}=\rho_{i}(g)\cdot W_{i}.$ (8)
In particular, by considering the component with index $1\in G$ on both sides of Equation 8, we see that $W_{i}(g^{-1})=W_{i}(1)\ \rho_{i}(g)$. We wish to show that $\rho_{i}$ is a homomorphism. To this end, for all $g,h\in G$ it holds that
By the surjectivity hypothesis, the set $\{W_{i}(g)\}_{g\in G}$ generates $\textnormal{End}(V_{i})$ as a vector space. Therefore, Equation 9 implies that $\rho_{i}(gh)=\rho_{i}(g)\rho_{i}(h)$, as desired. Lastly, note that $\rho_{i}$ is irreducible by the sujectivity assumption. Indeed, a nontrivial linear subspace of $V_{i}$ fixed by $\rho_{i}(g)$ for all $g$ would be sent by $W_{i}(g)$ into a fixed proper subspace due to Equation 8. This contradicts the surjectivity of $W_{i}$. ∎
Note that the surjectivity assumption implies the constraint $\sum_{i}\textnormal{dim}(V_{i})^{2}\leq|G|$. As a consequence of the result above, the full Fourier transform arises with an additional orthogonality assumption.

### v3

@@ Abstract
keywords:
Invariant neural networks, harmonic analysis, group representations
Figure 1: Weights learned by a neural network trained for invariance to planar rotations resemble circular harmonics. Data from [Sanborn et al. (2023)].
@@ 1 Introduction and Related Work
Artificial neural networks trained on natural data exhibit a striking phenomenon: regardless of exact initialization, dataset, or training objective, models trained on the same data domain frequently converge to similar learned representations (Li et al., 2015). For example, the early layer weights of diverse image models tend to converge to Gabor filters and color-contrast detectors (Olah et al., 2020). Remarkably, many of these same features are observed in the visual cortex (Hubel and Wiesel, 1959; Hass and Horwitz, 2013; Willeke et al., 2023), suggesting a form of representational universality that transcends biological and artificial substrates. While such findings are empirically well-established (Räuker et al., 2023), the field lacks theoretical explanations.
@@
Spatially localized versions of canonical 2D Fourier basis functions, such as Gabor filters or wavelets, are perhaps the most frequently observed universal features. They commonly arise in the early layers of vision models – trained with efficient coding (Olshausen and Field, 1997; Bell and Sejnowski, 1997), classification (Olah et al., 2020), temporal coherence (Hurri and Hyvärinen, 2003), and next-step prediction (Fiquet and Simoncelli, 2023) objectives – as well as in the primary visual cortices of diverse mammals – including cats (Hubel and Wiesel, 1962), monkeys (Hubel and Wiesel, 1968), and mice (Dräger, 1975). Non-localized Fourier features have been observed in networks trained to solve tasks that permit cyclic wraparound – for example, modular arithmetic (Nanda et al., 2023), more general group compositions (Chughtai et al., 2023), or invariance to the group of cyclic translations (Sanborn et al., 2023). In the domain of spatial navigation, the so-called grid cells of the entorhinal cortex (Moser et al., 2008) display periodic firing patterns at different spatial frequencies as they build a map of space. Their response properties are naturally modeled with the harmonics of the twisted torus (Guanella et al., 2007; Orchard et al., 2013; Gardner et al., 2022). Similar features also emerge in artificial neural networks trained to solve spatial navigation tasks (Banino et al., 2018; Cueva and Wei, 2018; Sorscher et al., 2019; Dorrell et al., 2022; Sorscher et al., 2023; Schaeffer et al., 2023). The ubiquity of these features across diverse learning systems is both striking and unexplained.
@@
In this work, we provide a mathematical explanation for the emergence of Fourier features in learning systems such as neural networks. We argue that the mechanism responsible for this emergence is the downstream invariance of the learner to the action of a group of symmetries (e.g., planar translations or rotations). Since natural data typically possess symmetries, invariance is a fundamental bias that is injected both implicitly and sometimes explicitly into learning systems (Bronstein et al., 2021; Cohen and Welling, 2016; Kondor and Trivedi, 2018). Motivated by this, we derive theoretical guarantees for the presence of Fourier features in invariant learners that apply to a broad class of machine learning models.
@@
Our results rely on the inextricable link between harmonic analysis and group theory (Folland, 2016). The standard discrete Fourier transform is a special case of more general Fourier transforms on groups, which can be defined by replacing the standard basis of harmonics by irreducible unitary group representations. The latter are equivalent to the familiar definition for cyclic or, more generally, commutative groups, but are more involved for non-commutative ones. In order to accommodate both scenarios, we develop a general theory that applies, in principle, to arbitrary finite groups.
@@
This work represents an attempt to provide mathematical grounding for a general algebraic theory of representation learning, while addressing the universality hypothesis for neural networks (Olah et al., 2020; Moschella et al., 2022). A suite of earlier theoretical works (Isely et al., 2010; Hillar and Sommer, 2015; Garfinkle and Hillar, 2019) established such universality for sparse coding models (Olshausen and Field, 1997), deriving the conditions under which a network will recover the original bases that generate data through sparse linear combinations. In this case, the statistics of the data determine the uniqueness of the representation. Our findings, on the other hand, are purely algebraic, since they rely exclusively on the invariance properties of the learner. Given the centrality of invariance to many machine learning tasks, our theory encompasses a broad class of scenarios and neural network architectures, while providing a new perspective on classical neuroscience (Pitts and McCulloch, 1947; Hubel and Wiesel, 1962). As such, it sets a foundation for a learning theory of representations in artificial and biological neural systems, grounded in the mathematics of symmetry.
@@ 1.1 Overview of Results
Here, the term “harmonic” refers to an irreducible unitary representation of $G$. In particular, one-dimensional unitary representations correspond to homomorphisms with the unit circle $\textnormal{U}(1)\subseteq\mathbb{C}$, which is reminiscent of the classical definition via the imaginary exponential. However, non-commutative groups can have higher-dimensional irreducible representations, intuitively meaning that harmonics are valued in unitary matrices. In this case, the components of $W$ can be interpreted as capsules in the sense of Hinton et al. (2018); i.e., neural units processing matrix-valued signals.
Since harmonic analysis is naturally formalized over the complex numbers, we consider models with complex weights $W$, which fits into the broader program of complex-valued machine learning (Bassey et al., 2021; Trabelsi et al., 2017; Löwe et al., 2022). We show that the hypothesis on $\varphi$ in Theorem 1.1 is satisfied by several machine learning models from the literature. In particular, the theorem applies to the recently-introduced (Bi)Spectral Networks (Sanborn et al., 2023), to single fully-connected layers of McCulloch-Pitts neurons, and, to an extent, to traditional deep networks. As an additional contribution, we generalize Spectral Networks to non-commutative groups.
The group-theoretical Fourier transform encodes the entire group structure of $G$. Therefore, as a consequence of Theorem 1.1, the multiplication table of $G$ can be recovered from the weights $W$ of an invariant parametric function $\varphi$ – a fact empirically demonstrated by Sanborn et al. (2023). This addresses the question of symmetry discovery – an established machine learning problem aiming to recover the unknown group of symmetries of data with minimal supervision and prior knowledge (Rao and Ruderman, 1998; Sohl-Dickstein et al., 2010; Desai et al., 2022). Since the multiplication table is a discrete object, it is expected that the invariance constraint on $\varphi$ can be loosened while still recovering the group correctly. To this end, we prove the following.
[…]
If $\varphi(W,x)$ is “almost invariant” to $G$ according to certain functional bounds and the weights are “almost orthonormal”, then the multiplication table of $G$ can be recovered from $W$.
[…]
@@ 2 Mathematical Background
We begin by introducing the fundamental concepts from harmonic analysis and group theory used in this paper. For a complete treatment, we refer the reader to Steinberg (2012).
@@ 2.1 Groups and Actions
Definition 2.1.
A group is a set $G$ equipped with a multiplication map $G\times G\rightarrow G$ denoted by $(g,h)\mapsto gh$, an inversion map $G\rightarrow G$ denoted by $g\mapsto g^{-1}$, and a distinguished identity element $1\in G$ such that for all $g,h,k\in G$:
@@
Examples of groups include the permutations of a set and the general linear group $\textnormal{GL}(V)$ of invertible operators over a vector space $V$, both equipped with the usual composition and inversion of functions. A further example that will be relevant in this work is the unitary group $\textnormal{U}(V)\subseteq\textnormal{GL}(V)$ associated to a Hilbert space $V$, consisting of operators $U$ satisfying $UU^{\dagger}=I$, where $I$ is the identity matrix. [^2] Given a linear operator $U\colon H\rightarrow H^{\prime}$ between Hilbert spaces, we denote by $U^{\dagger}\colon H^{\prime}\rightarrow H$ its adjoint operator, defined by $\langle Ux,y\rangle=\langle x,U^{\dagger}y\rangle$ for all $x\in H$, $y\in H^{\prime}$. Groups satisfying $gh=hg$ for all $g,h\in G$ are deemed commutative. The idea of a space $\mathcal{X}$ having $G$ as a group of symmetries is abstracted by the notion of group action.
Definition 2.2.
@@ 2.2 Harmonic Analysis on Groups
Harmonic analysis on groups (Folland, 2016) generalizes standard harmonic analysis. We focus here on finite groups for simplicity, which are sufficient for practical applications. This avoids technicalities such as integrability conditions and continuity issues arising for infinite groups. We start by considering commutative groups and cover non-commutative ones in Section 2.3.
Let $G$ be a finite commutative group of order $|G|$. Denote by $\langle G\rangle=\mathbb{C}^{G}$ the free complex vector space generated by $G$. Intuitively, an element $x=(x_{g})_{g\in G}\in\langle G\rangle$ represents a complex-valued signal over $G$. The space $\langle G\rangle$ is endowed with the convolution product:
@@
Definition 2.3.
The dual $G^{\vee}$ of $G$ is the set of homomorphisms $\rho:\ G\rightarrow\textnormal{U}(\mathbb{C})$, where $\textnormal{U}(\mathbb{C})\subseteq\mathbb{C}$ is the group of unitary complex numbers equipped with multiplication. It is itself a group when equipped with pointwise composition $(\rho\mu)(g)=\rho(g)\mu(g)$.
A homomorphism $\rho\in G^{\vee}$ intuitively represents a harmonic over $G$, generalizing the familiar notion from signal processing. If we endow $\langle G\rangle$ with the canonical scalar product $\langle x,y\rangle=\sum_{g\in G}\overline{x_{g}}y_{g}$, then $G^{\vee}\subseteq\langle G\rangle$ forms an orthogonal basis with all the norms equal to $\sqrt{|G|}$. The linear base-change is, by definition, the Fourier transform over $\langle G\rangle$:
Definition 2.4.
The Fourier transform is the map $\langle G\rangle\rightarrow\langle G^{\vee}\rangle$, $x\mapsto\hat{x}$, defined for $\rho\in G^{\vee}$ as $\hat{x}_{\rho}=\langle\rho,x\rangle$.
The Fourier transform is a linear isometry or, equivalently, a unitary operator, up to a multiplicative constant of $|G|$. Moreover, it exchanges the convolution product $\star$ over $\langle G\rangle$ with the Hadamard product $\odot$ over $\langle G^{\vee}\rangle$. Definition 2.4 generalizes the usual discrete Fourier transform in the following sense. For an integer $d>0$, consider the cyclc group $C_{d}$ with $d$ elements. Concretely, $C_{d}=\mathbb{Z}/d\mathbb{Z}$ is the group of integers modulo $d$ equipped with addition as composition. The dual $G^{\vee}$ consists of homomorphisms of the form $\mathbb{Z}/d\mathbb{Z}\ni g\mapsto e^{2\pi\sqrt{-1}gk/d}$ for $k\in\{0,\cdots,d-1\}$. Definition 2.4 specializes then to the familiar Fourier transform.
@@ 2.3 Non-Commutative Harmonic Analysis
So far, we have assumed that $G$ is commutative. In this section we briefly discuss the extension of Fourier theory to non-commutative groups. This however requires more elaborate theoretical tools, which we now introduce. To begin, in order to perform harmonic analysis on general groups it is necessary to discuss unitary representations. The latter will play the role of matrix-valued harmonics.
Definition 2.5.
@@
This generalizes Definition 2.4 since for a commutative group, $\rho_{V}$ is irreducible if, and only if, $\textnormal{dim}(V)=1$. Analogously to the commutative setting, the Fourier transform exchanges the convolution product $\star$ with the point-wise operator composition, which we still denote by $\odot$. Moreover, the Fourier transform is a unitary operator up to a multiplicative constant of $|G|$ with respect to the normalized Hilbert-Schmidt scalar product on $\textnormal{End}(V)$, given by $\langle A,B\rangle=\textnormal{dim}(V)\ \textnormal{tr}(A^{\dagger}B)$. The norm associated to the Hilbert-Schmidt scalar product is the Frobenius norm. The relations between irreducible unitary representations coming from the unitarity of the Fourier transform are known as Schur orthogonality relations.
@@ 3 Theoretical Results
Let $G$ be a finite group, $\mathcal{H}$ be a set, and $V_{1},\ldots,V_{k}$ be complex finite-dimensional Hilbert spaces. In what follows, we will consider the space:
@@
$\mathcal{W}$ is a Hilbert space when endowed with the scalar product given by the product of the canonical scalar product over $\langle G\rangle$ and the normalized Hilbert-Schmidt scalar products over $\textnormal{End}(V_{i})$. For $W\in\mathcal{W}$, we will denote each of its components as $W_{i}=(W_{i}(g))_{g}\in\textnormal{End}(V_{i})^{\oplus G}$. Moreover, we will often interpret elements $W\in\mathcal{W}$ as linear maps $\langle G\rangle\rightarrow\bigoplus_{i}\textnormal{End}(V_{i})$ via $W(x)=\sum_{g\in G}W(g)x_{g}$ for $x\in\langle G\rangle$, where $W(g)=(W_{i}(g))_{i}$. Note that $G$ acts on the left tensor factor of $\mathcal{W}$ while for every $i$, $\textnormal{U}(V_{i})$ acts on the right tensor factor of $\langle G\rangle\otimes\textnormal{End}(V_{i})$ by composition of operators.
Example 3.1.
Concretely, given coordinates on $V_{i}\simeq\mathbb{C}^{d_{i}}$ for $i=1,\ldots,k$, an element of $\mathcal{W}$ consists of a $k$-tuple of tensors $W_{i}\in\mathbb{C}^{d_{i}\times d_{i}\times|G|}\simeq\textnormal{End}(V_{i})^{\oplus G}$. For example, the dihedral group $D_{n}$ of isometries of a regular $n$-gon has $2n$ elements. For $n$ odd, it has $2$ irreducible unitary representations of dimension $1$, and $(n-1)/2$ of dimension $2$, implying that $\mathcal{W}$ consists of a direct sum of copies of $\mathbb{C}^{2n}$ and $\mathbb{C}^{2\times 2\times 2n}$.
Definition 3.2.
We say that a map $\varphi\colon\mathcal{W}\rightarrow\mathcal{H}$ has unitary symmetries if for all $W,W^{\prime}\in\mathcal{W}$ such hat $\|W_{i}\|=\|W_{i}^{\prime}\|$ for all $i$ and $\varphi(W)=\varphi(W^{\prime})$, we have that for every $i$ there exists a unitary operator $U_{i}\in\textnormal{U}(V_{i})$ such that $W_{i}=U_{i}\cdot W_{i}^{\prime}$.
In the context of machine learning, $\mathcal{H}$ will represent the hypothesis space, consisting of functions the model can learn. On the other hand, $\varphi$ will represent the parametrization of such hypotheses, with its domain $\mathcal{W}$ being the space of weights. Each component $\textnormal{End}(V_{i})$ of $\mathcal{W}$ will be responsible for parametrizing a computational unit, i.e. a complex-valued neuron in the language of neural networks. For commutative groups, we simply have $V_{i}=\mathbb{C}\simeq\textnormal{End}(V_{i})$. In general, $\textnormal{End}(V_{i})$ can be thought of as parametrizing matrix-valued signals, which, as mentioned in Section 1.1, are computed by neural units sometimes referred to as capsules (Hinton et al., 2018; Sabour et al., 2017). Lastly, the components of $\langle G\rangle$ will represent the input space. The fact that the latter consists of scalar signals over $G$ is a simplification of several practical scenarios. However, the results of this section can be generalized, to an extent, to signals over a set acted upon by $G$ – see Section 3.1. The following is an algebraic principle at the core of this work.
@@
  • $W_{i}$, seen as a linear map $\langle G\rangle\rightarrow\textnormal{End}(V_{i})$, is surjective for all $i$.
@@
We refer to the Appendix for a proof. Note that the surjectivity assumption implies the constraint $\textnormal{dim}(V_{i})^{2}\leq|G|$ for all $i$. As a consequence of the result above, the full Fourier transform arises with an additional orthogonality assumption.